the difference between a number and 15 is 7. find the number.
Found 2 solutions by Edwin McCravy, Alan3354:
Answer by Edwin McCravy(20056)

the difference between a number and 15 is 7. find the number.
x - 15 = 7

Add +15 to both sides:

x - 15 =   7
  + 15   +15
------------
x      =  21

Edwin

Answer by Alan3354(69443)

Edwin is close. I won't spend the time to make a big red font.
----------------------------------------------------------------
the difference between a number and 15 is 7. find the number.
------------
n - 15 = 7
n = 22
-------------
15-n = 7
n = 8